OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6 i think its 14 but im not sure .-.

jimthompson5910 (jim_thompson5910):

that's incorrect

OpenStudy (ja1):

First off distribute out the terms.

OpenStudy (ja1):

\[4x − 4 = 2x + 8 + 6 \] Right? @jim_thompson5910

OpenStudy (anonymous):

wait i just got soooo confused jim can you help me

OpenStudy (ja1):

Did you distribute the terms and get the equation above?

jimthompson5910 (jim_thompson5910):

4(x − 1) = 2(x + 2) + 6 4x − 4 = 2x + 4 + 6 ... distribute 4x - 4 = 2x + 10 4x - 2x = 10 + 4 ... ... ... x = ??

jimthompson5910 (jim_thompson5910):

btw JA1, you are close, but 2(x+2) turns into 2x+4 and not 2x+8

OpenStudy (ja1):

Why did I think 2 + 2 = 8 -.-

OpenStudy (ja1):

Sorry brain lag there I guess

OpenStudy (anonymous):

Lol uhm x = 8 (/.\)

jimthompson5910 (jim_thompson5910):

4x - 2x = 10 + 4 2x = 14 x = ??

jimthompson5910 (jim_thompson5910):

so I can see how you got 14 (somewhat)

OpenStudy (ja1):

You are close

OpenStudy (anonymous):

7 .-.

OpenStudy (ja1):

*applauds* :D good work

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

oh thank god thanks guys

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (ja1):

No prob that's what OpenStudy is all about, and welocme to OpenStudy if you have not yet been greeted :)

OpenStudy (anonymous):

Thanks JA1 C:

OpenStudy (ja1):

No problem, happy Open Studying and here you o for reference: http://openstudy.com/code-of-conduct